THE A-X TRANSITION IN Ar.NO AND PROGRESS IN THE ZEKE SPECTROSCOPY OF Ar.NO VIA THIS ELECTRONIC STATE

\underline{TIMOTHY G. WRIGHT}, D. M. SMITH, J. M. DYKE AND P. MACK, CHEMISTRY DEPART-MENT, UNIVERISTY OF SOUTHAMPTON, HIGHFIELD, SOUTHAMPTON, SO17 1BJ.

Recent work in our laboratory has concentrated upon the study of the Ar.NO complex using resonance-enhanced multiphoton ionization (REMPI) spectrosocopy. The spectrum of the A-X transition, although simple in appearance, is difficult to interpret; progress in this interpretation will be given. The transitions observed in the REMPI spectrum have been used in order to record zero-kinetic energy (ZEKE) photoelectron specra, and recent results on these experiments will be given.